Mass(100+-1)
Vol(20.0+-0.1)
Find density
1
Expert's answer
2019-01-22T08:41:07-0500

ρ=100/20.0=5.00

Δρ/ρ=√((1/100)^2+(0.1/20.0)^2 )=0.01118

Δρ=(0.01118)5.00= 0.06.

Density:

5.00±0.06